# DMS*: Minimizing Makespan for Multi-Agent Combinatorial Path Finding

### v1

@@ IV-A2 Binary Vector
For any joint vertex $v\in\mathcal{V^{W}}$, there can be multiple joint paths, e.g. $\pi_{1}(v_{o},v),\pi_{2}(v_{o},v)$, from $v_{o}$ to $v$ with different sets of targets visited along the joint path, and we need to differentiate between them during the search. We therefore introduce the following concepts. First, without losing generality, let all targets in $V_{t}$ be arranged as an ordered list $V_{t}=\{t_{m},m=1,2,\cdots,M\}$, where subscript $m$ indicates the index of a target in this list. [^3] In this paper, for a vector related to targets (e.g. a binary vector $\vec{a}$ of length $M$), we use subscripts (e.g. $a_{m}$) to indicate the elements in the vector. For a vector or joint vertex that is related to agents (e.g. $\vec{g}$ of length $N$, or $v\in\mathcal{V^{W}}$), we use a superscript (e.g. $i$ in $g^{i},v^{i}$) to indicate the element in the vector or joint vertex corresponding to an agent. Let $\vec{a}\in\{0,1\}^{M}$ denote a binary vector of length $M$ that indicates the visiting status of all targets in $V_{t}$, where the $m$-th component of $\vec{a}$ is denoted as $a_{m}$, and $a_{m}=1$ if the corresponding target $t_{m}$ is visited by some agent, and $a_{m}=0$ otherwise.
@@ IV-A3 Label
Let $l=(v,\vec{a},\vec{g})$ denote a label, where $v$ is a joint vertex, $\vec{a}$ is a binary vector of length $M$, and $\vec{g}$ is a cost vector of length $N$. Here, each component $g^{i},i\in I$ corresponds to the path cost of agent $i$. Intuitively, each label $l$ identifies a joint path from $v_{o}$ to $v$ that visits a subset of targets as described by $\vec{a}$ and with path costs as specified by $\vec{g}$. To simplify the presentation, given a label $l$, let $v(l),\vec{a}(l),\vec{g}(l)$ denote the corresponding component in $l$, and let $v^{i}(l)$ denote the vertex of agent $i$ in $v(l)$, $i\in I$. Furthermore, let $g_{max}(l):=\max_{i\in I}\vec{g}(l)$ denote the maximum path cost over all agents in $\vec{g}(l)$. To solve the MCPF-max in Def. 1, the algorithm needs to find a label $l$, whose corresponding joint path leads all agents to visit all targets and eventually reach the goals, and $g_{max}(l)$ is the objective value to be minimized.
@@ IV-A4 Label Comparison
To compare two labels at the same joint vertex (i.e., to compare the joint paths represented by these two labels), we need to compare both their binary vectors as well as the path costs. We introduce the following dominance relationship to compare two binary vectors.
Definition 2 (Binary Dominance).
@@
Intuitively, $\vec{a}$ dominates $\vec{b}$ if $\vec{a}$ visits all targets that are visited in $\vec{b}$, and $\vec{a}$ visits at least one more target than $\vec{b}$. In addition, two binary vectors are equal to each other ($\vec{a}=\vec{b}$) if both vectors are component-wise same to each other.
To compare two labels with the same joint vertex, we define the following dominance rule between them.
Definition 3 (Label Dominance).
For any two labels ${l}_{1}$ and ${l}_{2}$ with $v(l_{1})=v(l_{2})$, $l_{1}$ dominates $l_{2}$ ($l_{1}\succeq_{l}l_{2}$) if either of the following two conditions holds: (i) $\vec{a}(l_{1})\succeq_{b}\vec{a}(l_{2})$, $g_{max}(l_{1})\leq g_{max}(l_{2})$; (ii) $\vec{a}(l_{1})=\vec{a}(l_{2})$, $g_{max}(l_{1})<g_{max}(l_{2})$.
Intuitively, if $l_{1}$ dominates $l_{2}$, then the joint path identified by $l_{1}$ is guaranteed to be better than the joint path identified by $l_{2}$. If $l_{1}$ does not dominate $l_{2}$, $l_{1}$ is then non-dominated by $l_{2}$. Any two labels are non-dominated (with respect to each other) if each of them is non-dominated by the other. Two labels are said to be equal to (or same to) each other (notationally $l_{1}=l_{2}$) if they have the same vertex, binary vector and $g_{max}$ (i.e., $v(l_{1})=v(l_{2}),a(l_{1})=a(l_{2}),g_{max}(l_{1})=g_{max}(l_{2})$). Note that there is no need to compare $\vec{g}(l_{1})$ and $\vec{g}(l_{2})$ when comparing labels $l_{1},l_{2}$ since the problem in Def. III seeks to minimize the maximum path cost over the agents. Finally, for each joint vertex $v\in\mathcal{V^{W}}$, let $\mathcal{F}(v)$ denote a set of labels that are non-dominated to each other during the search.
@@ IV-A5 Target Sequencing
Let $\gamma^{i}=\{v^{i}_{o},t_{1},t_{2},\cdots,t_{k},v^{i}_{d}\}$ denote an (individual) target sequence, where $v^{i}_{o}\in V_{o}$ is the initial vertex of agent $i\in I$, each $t_{j}\in V_{t},j=1,2,\cdots,k$ is a target vertex and $v^{i}_{d}\in V_{d}$ is a goal vertex. A target sequence $\gamma^{i}$ specifies both the assignment and visiting order of targets for agent $i$. Let $\gamma=\{\gamma^{1},\gamma^{2},\cdots,\gamma^{N}\}$ denote a joint sequence, which specify the assignment and visiting order of all targets for all agents.
Given a label $l$, $\vec{a}(l)$ specifies the set of targets that are visited and unvisited. We introduce the notation $\gamma(l)=\{\gamma^{1}(l),\gamma^{2}(l),\cdots,\gamma^{N}(l)\}$, which is a joint sequence based on $l$ in the sense that $\gamma(l)$ visits all unvisited targets in $\vec{a}(l)$. In other words, (i) each $\gamma^{i}(l)\in\gamma(l)$ starts with $v^{i}(l)$, the vertex of agent $i$ in $v(l)$, visits a set of unvisited targets $\{t_{m}\}$, where the $m$-th component of $\vec{a}(l)$ is zero, and ends with a goal vertex $v^{i}_{d}$; (ii) all $\gamma^{i}(l),i\in I$ collectively visits all unvisited targets as specified in $\vec{a}(l)$.
@@ IV-A6 Target Graph
Let $\pi^{*}(u,v),u,v\in G^{W}$ denote a minimum-cost path between $u,v$ in the workspace graph $G^{W}$, and let $c_{\pi^{*}}(u,v)$ denote the cost of path $\pi^{*}(u,v)$. The cost of a target sequence $cost(\gamma^{i}(l))$ is equal to the sum of $c_{\pi^{*}}(u,v)$ for any two adjacent vertices $u,v$ in $\gamma^{i}(l)$.
Given a label $l$, to find a joint sequence $\gamma(l)$, a corresponding min-max multi-agent Hamiltonian path problem (mHPP) needs to be solved. Let SolveMHPP($l$) denote a procedure that solves the mHPP. SolveMHPP($l$) first formulates a target graph $G^{T}=(V^{T},E^{T},c^{T})$, which is computed based on the workspace graph $G^{W}$ as follows. The vertex set $V^{T}$ includes the current vertices of agents $v(l)$, the unvisited targets $\{t_{m}\in V_{t}|a_{m}(l)=0\}$ and goals $V_{d}$. $G^{T}$ is fully connected and $E^{T}=V^{T}\times V^{T}$. The edge cost in $G^{T}$ of any pair of vertices $u,v\in V^{T}$ is denoted as $c^{T}(u,v)$, which is the cost of a minimum cost path $\pi^{*}(u,v)$ in the workspace graph $G^{W}$. An example $G^{T}$ is shown in Fig. 2(b). The aforementioned target sequence $\gamma^{i}(l)$ for an agent $i$ is a path in $G^{T}$ and a joint sequence $\gamma(l)$ is a set of paths that starts from $v(l)$, visits all unvisited targets in $V_{t}$ as specified by $\vec{a}(l)$, and ends at goals $V_{d}$, while satisfying the assignment constraints. The procedure SolveMHPP$(l)$ can be implemented by various existing algorithms for mHPP.
@@ IV-A7 Heuristic and Policy
Given a joint sequence $\gamma(l)$, to simplify the notations, let $h^{i}(l):=cost(\gamma^{i}(l))$ be a heuristic value. Since agent-agent conflicts are ignored along the target sequences, the vector $\vec{h}:=\{h^{i}(l)|i\in I\}$ provides an estimate of the cost-to-go for each agent $i\in I$. When SolveMHPP($l$) solves the mHPP to optimality, the corresponding $\vec{h}$ provides lower bounds of the cost-to-go for all agents, which can be used as an admissible heuristic for the search. For any label $l$, let $\vec{f}(l):=\vec{g}(l)+\vec{h}(l)$ be the $f$-vector of $l$, which is an estimated cost vector of the entire joint path from the initial vertices to goals for all agents by further extending the joint path represented by $l$. Let $f_{max}(l):=\max_{i\in I}(f^{i}(l))$ denote the maximum component in $\vec{f}(l)$, which provides an estimate of the objective value related to label $l$.
Given a label $l$ and its joint sequence $\gamma(l)$, a joint policy $\phi_{\gamma(l)}$ can be built out of $\gamma(l)$, which maps one label to another label as follows. First, a joint path $\pi$ can be built based on $\gamma(l)$ by replacing any two subsequent vertices $u,v\in\gamma^{i}(l)$ for any agent $i\in I$ with a corresponding minimum cost path $\pi^{*}(u,v)$ in $G^{W}$. Then, along this joint path $\pi=(v_{0},v_{1},v_{2},\cdots,v_{\ell})$, all agents move from one joint vertex $v_{k}$ to $v_{k+1}$, $k=0,1,\cdots,\ell-1$. A set of corresponding binary vectors $\vec{a}_{k}$ for each $v_{k}\in\pi$ can be built by first making $\vec{a}_{0}=\vec{a}(l)$, and then updating $\vec{a}_{k},k=1,2,\cdots,\ell$ based on $\vec{a}_{k-1}$ and $v_{k}$ by considering if $v_{k}$ visits any new targets. A set of corresponding cost vectors $\vec{g}_{k}$ can also be computed in a similar way as $a_{k}$ by starting from $\vec{g}_{0}=\vec{g}(l)$. As a result, a joint policy $\phi_{\gamma(l)}$ is built by mapping one label $l_{k}=(v_{k},\vec{a}_{k},\vec{g}_{k})$ to the next label $l_{k+1}=(v_{k+1},\vec{a}_{k+1},\vec{g}_{k+1})$ along the target sequence $\phi_{\gamma(l)}$. Given a vertex $v^{i}$ of agent $i$, let $\phi^{i}_{\gamma(l)}(v^{i})$ denote the next vertex of agent $i$ in the joint policy $\phi_{\gamma(l)}$. An example of $\phi^{i}_{\gamma(l)}$ is shown in Fig. 2(d). A label $l$ is called on-policy if its next label is known in $\phi_{\gamma(l^{\prime})}$ for some $l^{\prime}$. Otherwise, $l$ is called off-policy, which means its next label is unknown yet and a mHPP needs to be solved from $l$ to find the policy $\phi_{\gamma(l)}$. We denote $\gamma(l),\vec{h}(l),\phi_{\gamma(l)}\leftarrow$SolveMHPP($l$) as the process of computing the joint sequence, heuristic values and joint policy.
@@ IV-B DMS* Algorithm
19:         if $I_{C}\neq\emptyset$ then
20:             continue
21:         if IsDominated($l^{\prime}$then
22:             DomBackProp($l,l^{\prime}$)
23:             continue
[…]
25:         $parent(l^{\prime})\leftarrow l$
26:         add $l^{\prime}$ to $\mathcal{F}(v(l^{\prime}))$ and back_set($l^{\prime}$)
27:         add $l^{\prime}$ OPEN with $f_{temp}(l^{\prime})$ as the priority
28:return Failure (no solution)
@@
1:INPUT: two labels $l,l^{\prime}$ where $l$ is the parent of $l^{\prime}$.
2:if $l^{\prime}$ is on-policy then
@@
4:    return
5:$\gamma(l^{\prime})\leftarrow\gamma(l)$
6:Compute $\phi_{\gamma(l^{\prime})}$ and $\vec{h}(l^{\prime})$ based on $\gamma(l^{\prime})$
7:if $\exists i\in I_{C}(l),g^{i}(l^{\prime})+h^{i}(l^{\prime})>f_{max}(l^{\prime})$ then
[…]
9:return
@@
3:    if $l\notin$ OPEN then
4:         add $l$ to OPEN
5:    for all $l^{\prime\prime}\in$ back_set($l$do
6:         BackProp($l^{\prime\prime}$, $I_{C}(l)$)
@@
To initialize (Lines 1-3), DMS* begins by creating an initial label $l_{o}$ and invokes the procedure SolveMHPP (as aforementioned) to compute $\gamma(l_{o}),\vec{h}(l_{o}),\phi_{\gamma(l_{o})}$ for $l_{o}$. For each label $l$, DMS* uses two $f$-values: $f_{temp}(l)$ and $f_{max}(l)$. Here, $f_{temp}(l)$ is a fast-to-compute yet roughly estimated cost-to-go, which does not require computing any joint sequence from $l$ to the goals. In contrast, $f_{max}(l)$ is an estimated cost-to-go based on a joint sequence, which is computationally more expensive to obtain than $f_{temp}(l)$. Similarly to A* [30], let OPEN denote a priority queue that stores labels $l$ and prioritizes them based on their $f$-values from the minimum to the maximum. In Alg. 1, we point out which $f$-value (either $f_{temp}$ or $f_{max}$) is used when a label is added to OPEN. Finally, $l_{o}$ is added to $\mathcal{F}(v_{o})$ since $l_{o}$ is non-dominated by any other labels at $v_{o}$, and $l_{o}$ is added to OPEN for future search. [^4] The notion of frontier set in this paper differs from the ones in . In , $\mathcal{F}(v)$ contains only expanded labels, which resembles the closed set in conventional A*, while in this paper, $\mathcal{F}(v)v\in\mathcal{V^{W}}$ includes both the labels that are expanded (closed labels) and labels that are generated and not yet expanded (open labels).
@@
In each search iteration (Lines 5-27), DMS* begins by popping a label $l$ from OPEN for processing. DMS* first calls the TargetSeq procedure for $l$, which takes the parent label of $l$ (denoted as $parent(l)$), $l$ itself, and the conflict set of $parent(l)$. TargetSeq determines either to call SolveMHPP to find a joint sequence for $l$ (i.e., $\gamma(l)$), or to re-use the joint sequence $\gamma(parent(l))$ that is previously computed for $parent(l)$ as $\gamma(l)$. We elaborate TargetSeq in Sec. IV-C. After TargetSeq on Line 7, the heuristic $\vec{h}(l)$ may change since a joint sequence may be computed for $l$ within TargetSeq, DMS* thus computes the $f_{max}(l)$ and compare it against $f_{temp}(l)$. When computing $f_{max}$ out of $\vec{g}$ and $\vec{h}$ on Line 8, a heuristic inflation factor $w\in[1,\infty)$ is used, which scales each component in $\vec{h}$ by the factor $w$. Heuristic inflation is a common technique for A* [31] and M*-based algorithms [10] that can often expedite the computation in practice while providing a $w$-bounded sub-optimal solution [31]. If $f_{max}(l)>f_{temp}(l)$, then $l$ should not be expanded in the current iteration since there can be labels in OPEN that have smaller $f$-value than $f_{max}(l)$. DMS* thus updates $f_{temp}(l)$ to be $f_{max}(l)$, re-adds $l$ to OPEN with the updated $f_{temp}(l)$, and skip the current search iteration. In a future search iteration, when this label $l$ is popped again, the condition on Line 9 will not hold since $f_{temp}(l)=f_{max}(l)$, and $l$ will be expanded.
@@
Afterwards, DMS* checks if $l$ leads to a solution using CheckSuccess($l$), which verifies if every component in $\vec{a}(l)$ is one and if every component of $v(l)$ is a unique goal vertex while satisfying the assignment constraints. If CheckSuccess($l$) returns true, a solution joint path $\pi^{*}$ is found and can be reconstructed by iterative tracking the parent pointers of labels from $l$ to $l_{o}$ in Reconstruct($l$). DMS* then terminates.
If CheckSuccess($l$) returns false, $l$ is expanded by considering its limited neighbors [10] described as follows. Let $I_{C}(l)\subseteq I$ denote the conflict set of label $l$, which is a subset of agents that are (or will be) in conflict. The limited neighbors of $l$ is a set of successor labels of $l$. For each agent $i$, if $i\notin I_{C}(l)$, agent $i$ is only allowed to move to its next vertex $\phi_{\gamma(l)}^{i}(v^{i}(l))$ as defined in the joint policy $\phi_{\gamma(l)}$. If $i\in I_{C}(l)$, agent $i$ is allowed to visit any adjacent vertex of $v^{i}(l)$ in $G^{W}$. In other words, the successor vertices of $v^{i}(l)$ are the following:
@@
Let $V_{succ}^{i}$ denote the set of successor vertices of $v^{i}(l)$, which is either of size one or equal to the number of edges incident on $v^{i}(l)$ in $G^{W}$. The successor joint vertices $V_{succ}$ of $v(l)$ is then the combination of $v^{i}(l)$ for all $i\in I$, i.e., $V_{succ}:=V_{succ}^{1}\times V_{succ}^{2}\times\cdots\times V_{succ}^{N}$. For each joint vertex $v^{\prime}\in V_{succ}$, a corresponding label $l^{\prime}$ is created and added to $L_{succ}$, the set of successor labels of $l$. When creating $l^{\prime}$, the corresponding $\vec{g}(l^{\prime})$ and $\vec{a}(l^{\prime})$ are computed based on $\vec{g}(l)$ $\vec{a}(l)$ and $v^{\prime}$.

### v2

@@ IV-A2 Binary Vector
For any $v\in\mathcal{V^{W}}$, there can be multiple joint paths, e.g. $\pi_{1}(v_{o},v),\pi_{2}(v_{o},v)$, from $v_{o}$ to $v$ with different sets of targets visited, and we need to differentiate between them. First, without losing generality, let all targets in $V_{t}$ be arranged as an ordered list $V_{t}=\{u_{m},m=1,2,\cdots,M\}$, where subscript $m$ indicates the index of a target in this list. [^3] In this paper, for a vector related to targets (e.g. a binary vector $\vec{a}$ of length $M$), we use subscripts (e.g. $a_{m}$) to indicate the elements in the vector. For a vector or joint vertex that is related to agents (e.g. $\vec{g}$ of length $N$, or $v\in\mathcal{V^{W}}$), we use a superscript (e.g. $i$ in $g^{i},v^{i}$) to indicate the element in the vector or joint vertex corresponding to an agent. Let $\vec{a}\in\{0,1\}^{M}$ denote a binary vector of length $M$ that indicates the visiting status of all targets in $V_{t}$ by a joint path during the search, where the $m$-th component of $\vec{a}$ is denoted as $a_{m}$, and $a_{m}=1$ if $u_{m}$ is visited, and $a_{m}=0$ otherwise.
@@ IV-A3 Label
Let $l=(v,\vec{a},\vec{g})$ denote a label, where $v$ is a joint vertex, $\vec{a}$ is a binary vector of length $M$, and $\vec{g}$ is a cost vector of length $N$. Here, each component $g^{i},i\in I$ is the path cost of agent $i$. During the search, each label $l$ identifies a joint path from $v_{o}$ to $v$ that visits a subset of targets as described by $\vec{a}$ and with path cost $\vec{g}$. Given $l$, let $v(l),\vec{a}(l),\vec{g}(l)$ denote the corresponding component in $l$, and let $v^{i}(l)$ denote the vertex of agent $i$ in $v(l)$, $i\in I$. Let $g_{max}(l):=\max_{i\in I}\vec{g}(l)$ denote the maximum path cost over all agents in $\vec{g}(l)$. To solve the MCPF-max in Def. 1, the planner searches for a label $l$, whose corresponding joint path leads all agents to visit all targets and eventually reaches the goals, and $g_{max}(l)$ is the objective value to be minimized.
@@ IV-A4 Label Comparison
To compare two labels at the same joint vertex, we compare both $\vec{a}$ and $\vec{g}$.
Definition 2 (Binary Dominance)
@@
Intuitively, $\vec{a}\succeq_{b}\vec{b}$ if $\vec{a}$ visits all targets that are visited in $\vec{b}$, and $\vec{a}$ visits at least one more target than $\vec{b}$. Two binary vectors are equal to each other ($\vec{a}=\vec{b}$) if both vectors are component-wise same to each other.
Definition 3 (Label Dominance)
For two labels ${l}_{1}$, ${l}_{2}$ with $v(l_{1})=v(l_{2})$, $l_{1}$ dominates $l_{2}$ ($l_{1}\succeq_{l}l_{2}$) if either (i) $\vec{a}(l_{1})\succeq_{b}\vec{a}(l_{2})$, $g_{max}(l_{1})\leq g_{max}(l_{2})$; or (ii) $\vec{a}(l_{1})=\vec{a}(l_{2})$, $g_{max}(l_{1})<g_{max}(l_{2})$ holds.
Intuitively, if $l_{1}\succeq_{l}l_{2}$, then the joint path identified by $l_{1}$ is guaranteed to be better than the joint path identified by $l_{2}$. If $l_{1}$ does not dominate $l_{2}$, $l_{2}$ is then non-dominated by $l_{1}$. Any two labels are non-dominated (with respect to each other) if each of them is non-dominated by the other. Two labels are said to be equal to (or same to) each other (notationally $l_{1}=l_{2}$) if $v(l_{1})=v(l_{2}),a(l_{1})=a(l_{2}),g_{max}(l_{1})=g_{max}(l_{2})$. There is no need to compare $\vec{g}(l_{1})$ and $\vec{g}(l_{2})$ when comparing labels $l_{1},l_{2}$ since the problem in Sec. III seeks to minimize the maximum path cost over the agents. Finally, for each joint vertex $v\in\mathcal{V^{W}}$, let $\mathcal{F}(v)$ denote a set of labels that are non-dominated to each other during the search.
@@ IV-A5 Target Sequencing
Let $\gamma^{i}=\{v^{i}_{o},u_{1},u_{2},\cdots,u_{k},v^{i}_{d}\}$ denote an (individual) target sequence, where each $u_{j},j=1,2,\cdots,k$ is a target vertex (i.e., $u_{j}\in V_{t}$). Let $\gamma=\{\gamma^{1},\gamma^{2},\cdots,\gamma^{N}\}$ denote a joint sequence, which specify the assignment and visiting order of all targets for all agents. Given $l$, $\vec{a}(l)$ specifies the set of targets that are visited and unvisited. We introduce the notation $\gamma(l)=\{\gamma^{1}(l),\gamma^{2}(l),\cdots,\gamma^{N}(l)\}$, a joint sequence based on $l$ in the sense that $\gamma(l)$ visits all unvisited targets in $\vec{a}(l)$: (i) each $\gamma^{i}(l)\in\gamma(l)$ starts with $v^{i}(l)$ ($v^{i}(l)$ is not necessarily an origin vertex), visits a set of unvisited targets $\{u_{m}\}$ (where the $m$-th component of $\vec{a}(l)$ is zero), and ends with a goal vertex $v^{i}_{d}$; (ii) all $\gamma^{i}(l),i\in I$ together visits all unvisited targets in $\vec{a}(l)$. Intuitively, $\gamma(l)$ is a joint sequence that is meant to complete the joint path represented by $l$.
@@ IV-A6 Target Graph
Let $\pi^{*}(u,v),u,v\in G^{W}$ denote a minimum-cost path between $u,v$ in $G^{W}$, and let $c_{\pi^{*}}(u,v)$ denote the cost of path $\pi^{*}(u,v)$. The cost of a target sequence $c(\gamma^{i}(l))$ is equal to the sum of $c_{\pi^{*}}(u,v)$ for any two adjacent vertices $u,v$ in $\gamma^{i}(l)$. Given $l$, to find $\gamma(l)$, a corresponding min-max multi-agent Hamiltonian path problem (mHPP) needs to be solved as follows. First, a target graph $G^{T}=(V^{T},E^{T},c^{T})$ is created based on $G^{W}$. The $V^{T}$ includes the current vertices of agents $v(l)$, the unvisited targets $\{u_{m}\in V_{t}|a_{m}(l)=0\}$ and goals $V_{d}$. $G^{T}$ is fully connected and $E^{T}=V^{T}\times V^{T}$. The edge cost in $G^{T}$ of any pair of $u,v\in V^{T}$ is denoted as $c^{T}(u,v)$, which is the cost of a minimum cost path $\pi^{*}(u,v)$ in the $G^{W}$. An example $G^{T}$ is shown in Fig. 2(b). A target sequence $\gamma^{i}(l)$ for an agent $i$ is a path in $G^{T}$, and a $\gamma(l)$ is a set of paths that starts from $v(l)$, visits all unvisited targets in $V_{t}$ as in $\vec{a}(l)$, and ends at goals $V_{d}$, while satisfying the assignment constraints. The procedure SolveMHPP$(l)$ can be implemented by various existing algorithms for mHPP.
@@ IV-A7 Heuristic and Policy
Given $\gamma(l)$, let $h^{i}(l):=c(\gamma^{i}(l))$ be a heuristic value. The vector $\vec{h}:=\{h^{i}(l)|i\in I\}$ provides an estimate of the cost-to-go for each agent $i\in I$. When SolveMHPP($l$) solves the mHPP to optimality, since conflicts are ignored along the target sequences, the corresponding $\vec{h}$ provides lower bounds of the cost-to-go for all agents, which is an admissible heuristic for the search. For any label $l$, let $\vec{f}(l):=\vec{g}(l)+\vec{h}(l)$ be the $f$-vector of $l$, which is an estimated cost vector of the entire joint path from $v_{o}$ to goals for all agents by further extending the joint path represented by $l$. Let $f_{max}(l):=\max_{i\in I}(f^{i}(l))$ denote the maximum component in $\vec{f}(l)$, which provides an estimate of the objective value related to $l$.
Given $l$ and $\gamma(l)$, a joint policy $\phi_{\gamma(l)}$ is built out of $\gamma(l)$, mapping one label to another as follows. First, a joint path $\pi$ is built based on $\gamma(l)$ by replacing any two subsequent vertices $u,v\in\gamma^{i}(l),i\in I$ with a corresponding minimum cost path $\pi^{*}(u,v)$ in $G^{W}$. Then, along this joint path $\pi=(v_{0},v_{1},v_{2},\cdots,v_{\ell})$, all agents move from $v_{k}$ to $v_{k+1}$, $k=0,1,\cdots,\ell-1$. The corresponding binary vectors $\vec{a}_{k}$ for each $v_{k}\in\pi$ are built by first making $\vec{a}_{0}=\vec{a}(l)$, and then updating $\vec{a}_{k},k=1,2,\cdots,\ell$ based on $\vec{a}_{k-1}$ and $v_{k}$ by checking if $v_{k}$ visits any new targets. The corresponding cost vectors $\vec{g}_{k}$ are computed similarly as $a_{k}$ by starting from $\vec{g}_{0}=\vec{g}(l)$. As a result, a joint policy $\phi_{\gamma(l)}$ is built by mapping one label $l=(v,\vec{a},\vec{g})$ to the next label $l^{\prime}=(v^{\prime},\vec{a}^{\prime},\vec{g}^{\prime})$ along the target sequence $\gamma(l)$. For a vertex $v^{i}$, let $\phi^{i}_{\gamma(l)}(v^{i})$ denote the next vertex of agent $i$ in the joint policy $\phi_{\gamma(l)}$. An example of $\phi^{i}_{\gamma(l)}$ is shown in Fig. 2(d). A label $l$ is on-policy if its next label is known in $\phi_{\gamma(l^{\prime})}$. Otherwise, $l$ is off-policy, i.e., the next label is unknown and a mHPP needs to be solved for $l$ to find the policy $\phi_{\gamma(l)}$. Let $\gamma(l),\vec{h}(l),\phi_{\gamma(l)}\leftarrow$SolveMHPP($l$) denote the process of computing the joint sequence, heuristic values and joint policy.
@@ IV-B DMS* Algorithm
19:         if $I_{C}\neq\emptyset$ continue then
20:         if IsDominated($l^{\prime}$then
21:             DomBackProp($l,l^{\prime}$)
22:             continue
23:         $f_{temp}(l^{\prime})\leftarrow\max_{i\in I}\{\vec{g}(l^{\prime})+w\cdot\textsl{SimpleHeu}(l^{\prime})\}$
24:         $parent(l^{\prime})\leftarrow l$
25:         add $l^{\prime}$ to $\mathcal{F}(v(l^{\prime}))$ and back_set($l^{\prime}$)
26:         add $l^{\prime}$ OPEN with $f_{temp}(l^{\prime})$ as the priority
27:return Failure (no solution)
@@
1:if $l^{\prime}$ is on-policy then
2:    $\gamma(l^{\prime})\leftarrow\gamma(l)$, return
@@
4:Compute $\phi_{\gamma(l^{\prime})}$ and $\vec{h}(l^{\prime})$ based on $\gamma(l^{\prime})$
5:if $\exists i\in I_{C}(l),g^{i}(l^{\prime})+h^{i}(l^{\prime})>f_{max}(l^{\prime})$ then
6:    $\gamma(l^{\prime}),\vec{h}(l^{\prime}),\phi_{\gamma(l^{\prime})}\leftarrow\textsl{SolveMHPP}(l^{\prime})$
7:return
@@
3:    if $l\notin$ OPEN then add $l$ to OPEN
4:    for all $l^{\prime\prime}\in$ back_set($l$do
5:         BackProp($l^{\prime\prime}$, $I_{C}(l)$)
@@
To initialize (Lines 1-3), DMS* creates an initial label $l_{o}$ and calls SolveMHPP to compute $\gamma(l_{o}),\vec{h}(l_{o}),\phi_{\gamma(l_{o})}$ for $l_{o}$. For each label $l$, DMS* uses two $f$-values: $f_{temp}(l)$ and $f_{max}(l)$, where $f_{temp}(l)$ is a fast-to-compute yet roughly estimated cost-to-go, which does not require computing any joint sequence from $l$ to the goals; and $f_{max}(l)$ is an estimated cost-to-go based on a joint sequence, which is computationally more expensive to obtain than $f_{temp}(l)$. Similarly to A* [21], let OPEN denote a priority queue storing labels and prioritizing labels based on their $f$-values from the minimum to the maximum. In Alg. 1, we point out which $f$-value (either $f_{temp}$ or $f_{max}$) is used when a label is added to OPEN. Finally, $l_{o}$ is added to $\mathcal{F}(v_{o})$ since $l_{o}$ is non-dominated by any other labels at $v_{o}$, and $l_{o}$ is added to OPEN for future search.
@@
In each iteration (Lines 5-26), DMS* pops a label $l$ from OPEN. DMS* calls TargetSeq for $l$, which takes the parent label of $l$ (denoted as $parent(l)$), $l$ itself, and the conflict set of $parent(l)$. TargetSeq either calls SolveMHPP to find a joint sequence $\gamma(l)$, or re-uses the joint sequence $\gamma(parent(l))$ that is previously computed for $parent(l)$. We elaborate TargetSeq in Sec. IV-C. After TargetSeq, $\vec{h}(l)$ may change since a joint sequence may be computed for $l$ within TargetSeq, DMS* thus computes $f_{max}(l)$ and compare it against $f_{temp}(l)$. When computing $f_{max}$ out of $\vec{g}$ and $\vec{h}$ on Line 8, a heuristic inflation factor $w\in[1,\infty)$ is used, which scales each component in $\vec{h}$ by the factor $w$. Heuristic inflation is common for A* [22] and M*-based algorithms [7] that can often expedite the computation in practice while providing a $w$-bounded sub-optimal solution [22]. If $f_{max}(l)>f_{temp}(l)$, then $l$ should not be expanded in the current iteration since there can be labels in OPEN that have smaller $f$-value than $f_{max}(l)$. DMS* thus updates $f_{temp}(l)$ to be $f_{max}(l)$, re-adds $l$ to OPEN with the updated $f_{temp}(l)$, and ends the iteration. In a future iteration, when this label $l$ is popped again, the condition on Line 9 will not hold since $f_{temp}(l)=f_{max}(l)$, and $l$ will be expanded.
@@
Afterwards, DMS* checks if $l$ leads to a solution using CheckSuccess($l$), which verifies if every component in $\vec{a}(l)$ is one and if every component of $v(l)$ is a unique goal vertex while satisfying the assignment constraints. If CheckSuccess($l$) returns true, a solution joint path $\pi^{*}$ is found and can be reconstructed by iterative tracking the parent pointers of labels from $l$ to $l_{o}$ in Reconstruct($l$). DMS* then terminates. If CheckSuccess($l$) returns false, $l$ is expanded by considering its limited neighbors [7] described as follows. The limited neighbors of $l$ is a set of successor labels of $l$. For each $i\in I$, if $i\notin I_{C}(l)$, agent $i$ is only allowed to move to its next vertex $\phi_{\gamma(l)}^{i}(v^{i}(l))$ as defined in the joint policy $\phi_{\gamma(l)}$. If $i\in I_{C}(l)$, agent $i$ is allowed to visit any adjacent vertex of $v^{i}(l)$ in $G^{W}$. The successor vertices of $v^{i}(l)$ are:
@@
Let $V_{succ}^{i}$ denote the set of successor vertices of $v^{i}(l)$, which is either of size one or equal to the number of edges incident on $v^{i}(l)$ in $G^{W}$. The successor joint vertices $V_{succ}$ of $v(l)$ is then the combination of $v^{i}(l)$ for all $i\in I$, i.e., $V_{succ}:=V_{succ}^{1}\times V_{succ}^{2}\times\cdots\times V_{succ}^{N}$. For each joint vertex $v^{\prime}\in V_{succ}$, a corresponding $l^{\prime}$ is created and added to $L_{succ}$, the set of successor labels of $l$. When creating $l^{\prime}$, the corresponding $\vec{g}(l^{\prime})$ and $\vec{a}(l^{\prime})$ are computed based on $\vec{g}(l)$ $\vec{a}(l)$ and $v^{\prime}$. In other words, the element in $\vec{g}(l^{\prime})$ is one unit larger than the corresponding element in $\vec{g}(l)$ (unless the agent has reached the goal and stays there) since every agents takes an action, and the element in $\vec{a}(l^{\prime})$ changes its value from 0 to 1, if $v^{\prime}$ visits any targets that are unvisited as in $a(l)$.